This triangle given below will be used to answer questions 4 - 6.
Use the Law of Sines to solve the triangle given the following information.
Round your answers to the nearest tenth for degrees and to the nearest hundredth for sides.
A = 60 degrees
a = 9
c = 10
C = __?___ Degrees

Answer by Boreal(15235)

a/sinA=c/sin C
9/sqrt(3)/2=10/sin C
sin C=18/sqrt(3)=0.96225
C=74.21 degrees